OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#9505 — Collect SFR in 06
Incident Report for Web Cloud
Resolved
We have an issue on the Collect SFR in Département 06 since 7am this morning.
A ticket has been opened. There's a cut in the fibre optic affecting multiple DSLAMs.
The intervention is underway at present.

Update(s):

Date: 2013-10-14 15:14:27 UTC
The fibre has been repaired. Access is back.
Posted Oct 14, 2013 - 10:34 UTC